OpenStudy (anonymous):

1 + 2x + 4x^2 + . . . = 3/4. What is the value of x, gives working out

OpenStudy (anonymous):

Are there infinite terms?

OpenStudy (raden):

use formula for solution left side : S=a/(1-r) a=the first term r=ratio

OpenStudy (anonymous):

a=1 And r=2x

OpenStudy (anonymous):

And S=3/4

OpenStudy (anonymous):

Now,can u do it?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i got all of that i just cant do it...

OpenStudy (anonymous):

i cannot reply sometimes cos it freezes

OpenStudy (anonymous):

U know the formula? S=a/(1-r)

OpenStudy (anonymous):

yes i know eveyrhitng.

OpenStudy (anonymous):

and i learn a(1-r)/1-r

Parth (parthkohli):

Wasn't it...\[\rm S = a{1 - r^n \over 1 - r}\]

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

i know everything i just need someone to do it so i can compare what i did

OpenStudy (anonymous):

i got something to the power of like 3/4 and then some weird fractrion

OpenStudy (anonymous):

Then substitute all the values and solve for x.

OpenStudy (anonymous):

i know but i didnt get it. who cares now. bye

Parth (parthkohli):

OK, if you got something, then tell us what you got.

OpenStudy (anonymous):

I mean the series converges .

OpenStudy (anonymous):

@Prudence are u there???